# Discrete Diffusion in Large Language and Multimodal Models: A Survey

### v3

@@ Abstract
In this work, we provide a systematic survey of Discrete Diffusion Language Models (dLLMs) and Discrete Diffusion Multimodal Language Models (dMLLMs). Unlike autoregressive (AR) models, dLLMs and dMLLMs adopt a multi-token, parallel decoding paradigm using full attention and a denoising-based generation strategy. This paradigm naturally enables parallel generation, fine-grained output controllability, and dynamic, response-aware perception. These capabilities are previously difficult to achieve with AR models. Recently, a growing number of industrial-scale proprietary d(M)LLMs, as well as a large number of open-source academic d(M)LLMs, have demonstrated performance comparable to their autoregressive counterparts, while achieving up to 10$\times$ acceleration in inference speed. These developments position discrete diffusion models as a promising alternative to intelligence based on the traditional autoregressive approach.
The advancement of discrete diffusion LLMs and MLLMs has been largely driven by progress in two domains. The first is the development of autoregressive LLMs and MLLMs, which has accumulated vast amounts of data, benchmarks, and foundational infrastructure for training and inference. A significant portion of these resources has been or can be transferred to support the development of dLLMs and dMLLMs. The second contributing domain is the evolution of the mathematical models underlying discrete diffusion. Originally based on the continuous-space, the diffusion models for LLM and MLLM have transitioned to the currently prevalent discrete-space diffusion models built on absorbing states. The adaptation and simplification of the mathematical framework has played a crucial role in enabling the scaling and engineering optimization of dLLMs and dMLLMs. Together, these advancements have catalyzed a surge in dLLMs and dMLLMs research in early 2025.
In this work, we present a comprehensive overview of the research in the dLLM and dMLLM domains. We trace the historical development of dLLMs and dMLLMs, formalize the underlying mathematical frameworks, and categorize representative models. We further analyze key techniques for training and inference, and summarize emerging applications across language, vision-language, and biological domains. We conclude by discussing future directions for research and deployment.
@@ I Introduction
Discrete Diffusion Large Language Models (dLLMs) and Discrete Diffusion Multimodal Large Language Models (dMLLMs)  [6, 7, 8, 9, 10, 11] have recently emerged as a promising direction. In contrast to AR generation, discrete diffusion models treat generation as an iterative denoising process over discrete token sequences. This paradigm eliminates the left-to-right constraint and enables generation that is parallel and structurally controllable with bidirectional attention mechanism.
@@
Originally proposed for continuous domains such as image generation [12], diffusion models have since been successfully extended to discrete spaces. Early efforts in this direction established the foundational mathematical formulations of discrete diffusion, introducing a token corruption schemes specifically designed for categorical data [13, 14]. These models demonstrated the feasibility of diffusion-based generation on various types of discrete data, including natural language [13, 15], images [13], and biological sequences such as DNA [16]. In this early stage, experiments were limited to models with around or fewer than 1 billion parameters. Through simplifications and reparameterizations [15, 16, 17], along with practical engineering efforts, the absorbing-state discrete diffusion formulation has gradually become the predominant mathematical framework adopted by open-source models. In this formulation, the inference process is an iterative masked token prediction, while the training process reduces to a masked language modeling task based on randomly sampled time indices and masked positions. The loss function can be written simply as a weighted cross-entropy loss. These simplifications significantly reduced the complexity and improved the stability in training and inference, paving the way for large-scale diffusion language models.
@@
Recent advances have significantly improved the scalability and effectiveness of discrete diffusion models [18, 19]. A major breakthrough on the industrial front came with the presentation of discrete diffusion-based large language models by Inception Labs and Google, namely Mercury [20] and Gemini Diffusion [21]. These models reports comparable performance on code and mathematics benchmarks with their AR counterpart, while also achieving 10$\times$ speedups in decoding, with about 1000 tokens per second.
@@
Alongside the development of open-source dLLMs and dMLLMs, significant progress has been made in optimizing training techniques specific to this paradigm. Strategies such as model initialization [18, 8] and masking scheduling [22] have been systematically explored to accelerate convergence, improve training stability, and enhance the final generation quality.
With the availability of open-source models, academic research has also deepened into the inference mechanisms of discrete diffusion. For example, during inference, the fundamental question of discrete diffusion models lies in determining, at each generation step, which tokens should be unmasked (i.e., finalized) and which should be left masked to future iterations. To this end, various unmasking [7, 8] and remasking [23, 24] strategies have been proposed, aiming to balance generation quality, stability, and speed. Another example is that, although discrete diffusion models are inherently capable of parallel decoding, they suffer from increased per-step computational cost due to the use of full attention masks. As a result, without optimization, their actual inference speed cannot surpass that of AR models. A considerable amount of work has focused on adapting existing inference acceleration algorithms originally designed for AR models to dLLMs and dMLLMs, aiming to reduce extensive full attention computations and enhance inference efficiency [25, 26].
@@
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=black, rounded corners, align=left, minimum width=4em, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, edge=-, opacity=1.0, line width=1.pt, , where level=1text width=12.0em,font=,edge=line width=1.pt, where level=2text width=12em,font=,edge=line width=1.pt, [Discrete Diffusion (M)LLMs, fill=Gray!10, edge=Black!60 [Mathematical Formulations (Sec. II), fill=lan!19, edge=lan!40 [Discrete Diffusion Model & Transi-
@@
tion Matrix (Sec. II.A), fill=lan!12, edge=lan!40 [Discrete Diffusion for Binary Variable [27],
Discrete Diffusion for Categorical Variable [14],
D3PM [13], Roulette Diffusion [28], GIDD [29], leaf, text width=18em, fill=lan!12, edge=lan!40] ] [Reparameterized Discrete Diffusion
Model [15] (Sec. II.B), fill=lan!12, edge=lan!40 ] [Simplified Masked Diffusion
@@
Model (Sec. II.C), fill=lan!12, edge=lan!40 [MLDM [16], MD4 [17], leaf, text width=18em, fill=lan!12, edge=lan!40] ] [Continuous Time Discrete Denoising
Models [30] (Sec. II.D), fill=lan!12, edge=lan!40 ] [Concrete Score (Sec. II.E), fill=lan!12, edge=lan!40 [CSM [31], DWDSE [16], Categorical Ratio Matching
@@
Loss [32], RADD [22], TCSM [33], CEDD [28], leaf, text width=18em, fill=lan!12, edge=lan!40] ] [Discrete Flow Matching [24]
(Sec. II.F), fill=lan!12, edge=lan!40 ] [Block Diffusion Models [34]
@@
(Sec. II.G), fill=lan!12, edge=lan!40 ] ] [Representative Models (Sec. III), fill=mul!19, edge=mul!40 [Discrete Diffusion Models around
1B (Sec. III.A), fill=mul!12, edge=mul!40 [ D3PM [13], DiffusionBERT [35], RDM [15],
Masked-Diffuse LM [36], Diffusion-NAT [37],
TESS [38], Plaid [39], SEDD [40], MDLM [16],
@@
MD4 [17], UniDisc [41] , leaf, text width=18em, fill=mul!12, edge=mul!40] ] [Large Diffusion Language Models
(Sec. III.B), fill=mul!12, edge=mul!40 [ LLaDA [6], DIFFUSION-LLMs[42], DiffuGPT &
DiffuLLaMA[18], DREAM[7], LLaDA 1.5 [43],
@@
TESS 2[44] , leaf, text width=18em, fill=mul!12, edge=mul!40] ] [Large Diffusion Multimodal Models
@@
(Sec. III.C), fill=mul!12, edge=mul!40 [ Dimple [8], LaViDa [10], LLaDA-V [9] , leaf, text width=18em, fill=mul!12, edge=mul!40] ] [Large Unified Models (Sec. III.D), fill=mul!12, edge=mul!40 [ MMaDA [11], FUDOKI [45], Muddit [46] , leaf, text width=18em, fill=mul!12, edge=mul!40] ] ] [Training Techniques (Sec. IV), fill=uni!19, edge=uni!40 [Initialization Technique (Sec. IV.B), fill=uni!12, edge=uni!40 [ BERT Initialization [35], Autoregressive Initiali-
zation[18, 7], Autoregressive-then Diffusion
@@
Training [8] , leaf, text width=18em, fill=uni!12, edge=uni!40] ] [Complementary Masking Technique
[10] (Sec. IV.C), fill=uni!12, edge=uni!40 ] [Masking Scheduling Technique
(Sec. IV.D), fill=uni!12, edge=uni!40 [ Linear Schedule [13], Geometric Schedule [40], Cosine
@@
Schedule [47], Token-wise Masking Scheduling [35] , leaf, text width=18em, fill=uni!12, edge=uni!40] ] [Addressing Training-Testing
@@
Discrepancy (Sec. IV.E), fill=uni!12, edge=uni!40 [ Two-Step Strategy [48] , leaf, text width=18em, fill=uni!12, edge=uni!40] ] [Reweighting Technique (Sec. IV.F), fill=uni!12, edge=uni!40 [ MGDM [49] , leaf, text width=18em, fill=uni!12, edge=uni!40] ] [Distillation (Sec. IV.G), fill=uni!12, edge=uni!40 [ Di4C [50] , leaf, text width=18em, fill=uni!12, edge=uni!40] ] ] [Inference Techniques (Sec. V), fill=SkyBlue!19, edge=SkyBlue!40 [Unmasking Techniques (Sec. V.A), fill=SkyBlue!12, edge=SkyBlue!40 [ Metrics Used in Unmasking: Confidence, Margin,
Entropy [7], Confident Decoding [8],
Semi-Autoregressive Unmasking [9],
@@
Continuous Time (Flow Matching) Unmasking [24] , leaf, text width=18em, fill=SkyBlue!12, edge=SkyBlue!40] ] [Remasking Techniques (Sec. V.B), fill=SkyBlue!12, edge=SkyBlue!40 [ Discrete Time Remasking [23], Continuous Time
@@
(Flow Matching) Remasking [24] , leaf, text width=18em, fill=SkyBlue!12, edge=SkyBlue!40] ] [Prefilling and Caching Technique
(Sec. V.C), fill=SkyBlue!12, edge=SkyBlue!40 [ Prefilling [8, 10], dKV-Cache [25], dLLM-Cache [26],
@@
DualCache [51] , leaf, text width=18em, fill=SkyBlue!12, edge=SkyBlue!40] ] [Guidance Technique (Sec. V.D), fill=SkyBlue!12, edge=SkyBlue!40 [ Classifier-Free Guidance [19], Classifier Guidance [52],
Reward Guidance [44], EDLM [53] , leaf, text width=18em, fill=SkyBlue!12, edge=SkyBlue!40] ] ] [Applications (Sec. VI), fill=YellowOrange!19, edge=YellowOrange!40 [Text Generation and Style Control
(Sec. VI.A), fill=YellowOrange!12, edge=YellowOrange!40 [ DiffuSeq_StylePTB [54], DiffEmbed [55], SLD [56],
DiffusPoll [57], PoetryDiffusion [58] , leaf, text width=18em, fill=YellowOrange!12, edge=YellowOrange!40] ] [Text Editing and Summarization
(Sec. VI.B), fill=YellowOrange!12, edge=YellowOrange!40 [ EdiText [59], CrossMamba [60], DiffETM [61],
TermDiffuSum [62] , leaf, text width=18em, fill=YellowOrange!12, edge=YellowOrange!40] ] [Sentiment Analysis and Data
@@
Augmentation (Sec. VI.C), fill=YellowOrange!12, edge=YellowOrange!40 [ $\textrm{CDA}^{2}$ [63], DiffusionCLS [64], GDP [65] , leaf, text width=18em, fill=YellowOrange!12, edge=YellowOrange!40] ] [Knowledge and Reasoning
(Sec. VI.D), fill=YellowOrange!12, edge=YellowOrange!40 [ DoT [66], DiffuCOMET [67], DPCL-Diff [68],
@@
d1 [69], MGDM [49] , leaf, text width=18em, fill=YellowOrange!12, edge=YellowOrange!40] ] [Vision and Multimodal (Sec. VI.E), fill=YellowOrange!12, edge=YellowOrange!40 [ DiffVLA [70], UDAN-CLIP [71], VPDD [72],
@@
M2D2M [73], AR-Diffusion [74] , leaf, text width=18em, fill=YellowOrange!12, edge=YellowOrange!40] ] [Biological and Drug Discovery
(Sec. VI.F), fill=YellowOrange!12, edge=YellowOrange!40 [ MolEditRL [75], CFP-Gen [76], TransDLM [77],
GenMol [78], DPLM-2 [79], PepTune [80] , leaf, text width=18em, fill=YellowOrange!12, edge=YellowOrange!40] ] ] ]
@@
Figure 2: Overview of our survey. We begin by introducing the mathematical foundations of discrete diffusion language models (Sec. II). Next, we present a high-level overview of representative base models (Sec. III), followed by discussions on training strategies (Sec. IV) and inference techniques (Sec. V). In addition, we also review a wide range of applications (Sec. VI) that adopt discrete diffusion language models as their core model.
@@
In the rest of this paper, Sec. II presents the mathematical foundations of discrete diffusion models. The core of mathematical formulations lies in the definitions of the forward and reverse diffusion processes over discrete spaces. These are modeled via Markov chains that incrementally corrupt and denoise sequences. The framework encompasses both discrete-time and continuous-time formulations, supporting extensions such as flow matching and semi-autoregressive block diffusion model.Sec. III surveys representative discrete diffusion language models across varying scales. This includes early-stage models (e.g., D3PM [13], RDM [15], DiffusionBERT [35]), scaled dLLMs [6, 18, 7],dMLLMs [8, 10, 9], and unified models [11]. Sec. IV discusses the key training strategies used in dLLMs and dMLLMs, including initialization, masking strategies, and masking schedules. Sec. V lists various inference techniques used in dLLMs and dMLLMs, including unmasking, remasking, caching and guidance. Sec. VI reviews the broad range of applications powered by discrete diffusion models. Finally, Sec. VII summarize potential directions for future research. The organization of this survey is illustrated in Fig. 2.

### v4

@@ Abstract
[…]
@@ I Introduction
Discrete Diffusion Large Language Models (dLLMs) and Discrete Diffusion Multimodal Large Language Models (dMLLMs)  [6, 7, 8, 9, 10, 11] have recently emerged as a promising direction. In tasks such as code generation [12, 13], planning [7], and Sudoku [7], dLLMs have been widely shown to achieve better performance than autoregressive models. Moreover, [14] demonstrates that in a data-constrained setting, increasing the number of training FLOPs for dLLMs enables it to outperform AR.
In contrast to AR generation, discrete diffusion models treat generation as an iterative denoising process over discrete token sequences. This paradigm eliminates the left-to-right constraint and enables generation that is parallel and structurally controllable with bidirectional attention mechanism.
@@
Originally proposed for continuous domains such as image generation [15], diffusion models have since been successfully extended to discrete spaces. Early efforts in this direction established the foundational mathematical formulations of discrete diffusion, introducing a token corruption schemes specifically designed for categorical data [16, 17]. These models demonstrated the feasibility of diffusion-based generation on various types of discrete data, including natural language [16, 18], images [16], and biological sequences such as DNA [19]. In this early stage, experiments were limited to models with around or fewer than 1 billion parameters. Through simplifications and reparameterizations [18, 19, 20], along with practical engineering efforts, the absorbing-state discrete diffusion formulation has gradually become the predominant mathematical framework adopted by open-source models and is termed as the masked diffusion model.
@@
[…]
@@
Alongside the development of open-source dLLMs and dMLLMs, significant progress has been made in optimizing training and decoding techniques specific to this paradigm. For example, strategies such as masking scheduling [23] and reinforcement learning [24, 11] have been systematically explored to accelerate convergence, improve training stability, and enhance the final generation quality; various unmasking [7, 8], remasking [25, 26], and sampling [27, 28], and caching [29, 30] strategies have been proposed to balance generation quality, stability, and speed.
@@
for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=left, font=, rectangle, draw=black, rounded corners, align=left, minimum width=4em, s sep=3pt, inner xsep=2pt, inner ysep=3pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, edge=-, opacity=1.0, line width=1.pt, , where level=1text width=13em,font=,edge=line width=1.pt, where level=2text width=12em,font=,edge=line width=1.pt, where level=3text width=20em,font=,edge=line width=1.pt, [Discrete Diffusion (M)LLMs, fill=Gray!10, edge=Black!60 [Mathematical Formulations (Sec. II), fill=lan!19, edge=lan!40 [Discrete Diffusion Model & Transi-
@@
tion Matrix (Sec. II.A), fill=lan!12, edge=lan!40 [Discrete Diffusion for Binary [31] and Categorical [17] Vari-
@@
able, D3PM [16], Roulette Diffusion [32], GIDD [33], leaf, fill=lan!12, edge=lan!40] ] [Reparameterized Discrete Diffusion Model [18] (Sec. II.B), fill=lan!12, edge=lan!40, text width=20em ] [Simplified Masked Diffusion Model (Sec. II.C), fill=lan!12, edge=lan!40, text width=17em [MLDM [19], MD4 [20], leaf, fill=lan!12, edge=lan!40, text width=8em] ] [Continuous Time Discrete Denoising Models [34] (Sec. II.D), fill=lan!12, edge=lan!40, text width=21em ] [Concrete Score (Sec. II.E), fill=lan!12, edge=lan!40, text width=10em [CSM [35], DWDSE [19], Categorical Ratio Matching Loss [36],
@@
RADD [23], TCSM [37], CEDD [32], leaf, fill=lan!12, edge=lan!40, text width=22em] ] [Discrete Flow Matching [26] (Sec. II.F), fill=lan!12, edge=lan!40, text width=14em ] ] [Modeling Language Diffusion (Sec. III), fill=red!10, edge=red!30 [Block Diffusion Models [38](Sec. III.A), Flexible-Length Masked Diffusion [39](Sec. III.B),
@@
Partial Masking [40](Sec. III.C), DDOT [41](Sec. III.D), fill=red!10, edge=red!30, text width=33.6em ] ] [Representative Models (Sec. IV), fill=mul!19, edge=mul!40 [Discrete Diffusion Models around
1B (Sec. IV.A), fill=mul!12, edge=mul!40 [ D3PM [16], DiffusionBERT [42], RDM [18],Masked-Diffuse
LM [43], Diffusion-NAT [44], TESS [45], Plaid [46], SEDD
@@
[47], MDLM [19], MD4 [20], UniDisc [48] , leaf, fill=mul!12, edge=mul!40] ] [Large Diffusion Language Models
(Sec. IV.B), fill=mul!12, edge=mul!40 [ LLaDA [6],DIFFUSION-LLMs[49],DiffuGPT&DiffuLLaMA
[21], DREAM[7], LLaDA 1.5 [24], TESS 2[50], DreamOn
@@
[51], DreamCoder [52], DiffuCoder [53], Seed Diffusion [54] , leaf, fill=mul!12, edge=mul!40] ] [Large Diffusion Multimodal Models
@@
(Sec. IV.C), fill=mul!12, edge=mul!40 [ Dimple [8], LaViDa [10], LLaDA-V [9] , leaf, fill=mul!12, edge=mul!40] ] [Large Unified Models (Sec. IV.D), fill=mul!12, edge=mul!40 [ MMaDA [11], FUDOKI [55], Muddit [56] , leaf, fill=mul!12, edge=mul!40] ] ] [Training Techniques (Sec. V), fill=uni!19, edge=uni!40 [Initialization Technique (Sec. V.B), fill=uni!12, edge=uni!40 [ BERT Initialization [42], Autoregressive Initialization[21, 7],
@@
Autoregressive-then Diffusion Training [8] , leaf, fill=uni!12, edge=uni!40] ] [Complementary Masking[10](Sec. V.C), fill=uni!12, edge=uni!40, text width=13em ] [Masking Scheduling Technique
(Sec. V.D), fill=uni!12, edge=uni!40 [ Linear [16],Geometric [47, 20],Cosine [57],Polynomial [20],
@@
Token-wise Masking Scheduling [42], Blockwise SFT [58] , leaf, fill=uni!12, edge=uni!40] ] [Addressing Training-Testing
@@
Discrepancy (Sec. V.E), fill=uni!12, edge=uni!40 [ Two-Step Strategy [59] , leaf, fill=uni!12, edge=uni!40] ] [Reweighting Technique (Sec. V.F), fill=uni!12, edge=uni!40 [ MGDM [60] , leaf, fill=uni!12, edge=uni!40] ] [Distillation (Sec. V.G), fill=uni!12, edge=uni!40 [ Di4C [61] , leaf, fill=uni!12, edge=uni!40] ] [Reinforcement Learning (Sec. V.H), fill=uni!12, edge=uni!40 [ UniGRPO [11],VRPO [24],SDPO [62],wd1 [63],DCoLT [64] , leaf, fill=uni!12, edge=uni!40] ] ] [Inference Techniques (Sec. VI), fill=SkyBlue!19, edge=SkyBlue!40 [Unmasking Techniques (Sec. VI.A), fill=SkyBlue!12, edge=SkyBlue!40 [ Metrics Used in Unmasking: Confidence, Margin, Entropy
[7, 65], EB-Sampler [66], PC-Sampler. [67]; Confident
Decoding [8], Block-wise [9], DUS [68], Continuous Time
@@
(Flow Matching) Unmasking [26] , leaf, fill=SkyBlue!12, edge=SkyBlue!40] ] [Remasking Techniques (Sec. VI.B), fill=SkyBlue!12, edge=SkyBlue!40 [ Discrete Time Remasking [25],Wide In Narrow Out [69]
@@
Continuous Time (Flow Matching) Remasking [26] , leaf, fill=SkyBlue!12, edge=SkyBlue!40] ] [Prefilling and Caching Technique
(Sec. VI.C), fill=SkyBlue!12, edge=SkyBlue!40 [ Prefilling [8, 10], dKV-Cache [29], dLLM-Cache [30],
@@
DualCache [70] , leaf, fill=SkyBlue!12, edge=SkyBlue!40] ] [Guidance Technique (Sec. VI.D), fill=SkyBlue!12, edge=SkyBlue!40 [ Classifier-Free Guidance [22], Classifier Guidance [71],
Reward Guidance [50], EDLM [72] , leaf, fill=SkyBlue!12, edge=SkyBlue!40] ] [Sampling Technique (Sec. VI.E), fill=SkyBlue!12, edge=SkyBlue!40 [ Early Stopping [27, 28], Particle Gibbs Sampling [73],
Temporal Self-Consistency [74], leaf, fill=SkyBlue!12, edge=SkyBlue!40] ] [ Context Length Extension (Sec. VI.F), fill=SkyBlue!12, edge=SkyBlue!40 [ LongLLaDA [75], leaf, fill=SkyBlue!12, edge=SkyBlue!40 ] ] [Sparse Computation (Sec. VI.G), fill=SkyBlue!12, edge=SkyBlue!40 [ Sparse-dLLM [76], DPad [77],Pipelined Parallel [78], leaf, fill=SkyBlue!12, edge=SkyBlue!40] ] [Response Length Control (Sec. VI.H), fill=SkyBlue!12, edge=SkyBlue!40 [ DAEDAL [79], leaf, fill=SkyBlue!12, edge=SkyBlue!40] ] ] [Privacy and Safety (Sec. VII), fill=Goldenrod!19, edge=Goldenrod!40 [ DIJA [80], PAD [81], MOSA [82], fill=Goldenrod!12, edge=Goldenrod!40 ] ] [Quantization (Sec. VIII), fill=SpringGreen!19, edge=SpringGreen!40 [Quantization Meets dLLMs[83],
DLLMQuant [84], fill=SpringGreen!12, edge=SpringGreen!40 ] ] [Applications (Sec. IX), fill=YellowOrange!19, edge=YellowOrange!40 [Language Related Applications
(Sec. IX.A), fill=YellowOrange!12, edge=YellowOrange!40 [DiffuSeq_StylePTB [85], DiffEmbed [86], SLD [87],
DiffusPoll [88], PoetryDiffusion [89], SVDD [90],
EdiText [91], CrossMamba [92], DiffETM [93],
TermDiffuSum [94], $\textrm{CDA}^{2}$ [95], DiffusionCLS [96], GDP
@@
[97], Layout-Corrector [98] , leaf, fill=YellowOrange!12, edge=YellowOrange!40] ] [Knowledge and Reasoning
(Sec. IX.B), fill=YellowOrange!12, edge=YellowOrange!40 [ DoT [99], DiffuCOMET [100], DPCL-Diff [101],
@@
d1 [102], MGDM [60], NeSyDM [103], leaf, fill=YellowOrange!12, edge=YellowOrange!40] ] [Vision and Multimodal (Sec. IX.C), fill=YellowOrange!12, edge=YellowOrange!40 [ UDAN-CLIP [104], M2D2M [105], AR-Diffusion [106] , leaf, fill=YellowOrange!12, edge=YellowOrange!40] ] [Robotics and Autonomous Driving
(Sec. IX.D), fill=YellowOrange!12, edge=YellowOrange!40 [ DiffVLA [107], DDVLA [108], ViLaD [109], VPDD [110],
DGD [111] , leaf, fill=YellowOrange!12, edge=YellowOrange!40] ] [Graph and Structured Predictions
@@
(Sec. IX.E), fill=YellowOrange!12, edge=YellowOrange!40 [ LO-ARM [112], ReDiSC [113], Scaffold Diffusion [114] , leaf, fill=YellowOrange!12, edge=YellowOrange!40] ] [Biological and Drug Discovery
(Sec. IX.F), fill=YellowOrange!12, edge=YellowOrange!40 [ MolEditRL [115], CFP-Gen [116], TransDLM [117],
GenMol [118], DPLM-2 [119], PepTune [120],
CMCM-DLM [121] , leaf, fill=YellowOrange!12, edge=YellowOrange!40] ] ] ]
@@
Figure 2: Overview of our survey. We begin by introducing the mathematical foundations of discrete diffusion language models (Sec. II). Next, we present a high-level overview of representative base models (Sec. IV), followed by discussions on training strategies (Sec. V) and inference techniques (Sec. VI). In addition, we also review a wide range of applications (Sec. IX) that adopt discrete diffusion language models as their core model.
@@
In the rest of this paper, Sec. II presents the mathematical foundations of discrete diffusion models. The core of mathematical formulations lies in the definitions of the forward and reverse diffusion processes over discrete spaces. These are modeled via Markov chains that incrementally corrupt and denoise sequences. The framework encompasses both discrete-time and continuous-time formulations, supporting extensions such as flow matching and semi-autoregressive block diffusion model.Sec. IV surveys representative discrete diffusion language models across varying scales. This includes early-stage models (e.g., D3PM [16], RDM [18], DiffusionBERT [42]), scaled dLLMs [6, 21, 7], dMLLMs [8, 10, 9], and unified models [11]. Sec. V discusses the key training strategies used in dLLMs and dMLLMs, including initialization, masking strategies, and masking schedules. Sec. VI lists various inference techniques used in dLLMs and dMLLMs, including unmasking, remasking, caching and guidance. Sec. IX reviews the broad range of applications powered by discrete diffusion models. Finally, Sec. X summarize potential directions for future research. The organization of this survey is illustrated in Fig. 2.